OpenStudy (anonymous):

Find the zeros of f(x)=(x-4)^2(x+3)^3 and state the multiplicity.

OpenStudy (mathstudent55):

Similar to previous problem: (x - 4)^2(x + 3)^3 = 0 (x - 4)(x + 3) =0 Get solutions for x from step above. Then each answer has multiplicity of the exponent of its binomial.